A 10 kg mass needs to be lifted onto a truck that is 1.5 m from the ground. A ramp, 10 m long, was used to get the mass onto the truck. What is the efficiency of the move?


1
Expert's answer
2021-01-14T10:38:00-0500
"E=mgh=(10)(9.8)(1.5)=150\ J"

"\eta=\frac{E}{Fs}=\frac{150}{10F}=\frac{15}{F}"